# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks – the GATTACA Framework

### v2

@@ Abstract
Index Terms:
Boolean Networks, Network Control, Deep Reinforcement Learning, Gene Regulatory Networks, Cellular Reprogramming
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Complex diseases remain a major challenge for biomedical research due to the intricate interactions between genes forming gene regulatory networks (GRNs). These networks exhibit complex dynamics emerging from the structure of the network: the state of the genome evolves in time, finally reaching some stable gene expression profiles characterising cell types or, more generally, stable cellular functional states [1].
@@
In some cases, when a GRN is perturbed, it can reach an “unhealthy” state, that is otherwise inaccessible [2]. An artificial change of the cell fate, referred to as cellular reprogramming, aims to solve this issue by steering the GRN back to a “healthy” state. However, identifying effective reprogramming targets and intervention strategies through wet-lab experiments alone is prohibitively expensive and time-consuming. This motivates the development of in-silico methods.
@@
Capturing emergent properties of GRN dynamics requires the consideration of the system as a whole. To this end, GRNs can be holistically modelled with the formalism of Boolean networks (BNs) originally introduced by Kauffman in [3]. Its simplicity enables tractable simulations and analysis of non-linear biological dynamics while retaining essential, emergent complex system behaviour. Cellular reprogramming can be formulated as a control problem within the framework of BNs.
@@
To support cellular reprogramming, control strategies must be computed for large-scale BNs, taking into account both the structure and dynamics of the network [4]. Scalability is the main concern when devising new computational methods for control of asynchronous BNs. Existing state-of-the-art methods and tools, such as CABEAN [5], provide exact algorithms for attractor-based control. Nevertheless, they are limited in scalability because of the infamous “curse of dimensionality” problem, i.e. the size of the state space grows exponentially with the number of genes considered, which is handled by additional assumptions about network decomposability.
@@
Deep reinforcement learning (DRL) has proven highly successful in decision problems characterised by very large state-action spaces [6]. Therefore, it seems a promising alternative to traditional methods for controlling GRNs. In [7], pbn-STAC – a DRL-based tool for control of asynchronous BNs and probabilistic Boolean networks (PBNs) was introduced, which was a result of exploration into the potential of DRL for developing scalable control methods for both BN and PBN models of GRNs in the context of cellular reprogramming. Therein, notions of a pseudo-attractor and a pseudo-attractor state were introduced, a pseudo-attractor states identification procedure (PASIP) was devised, and a novel control problem, i.e. the Source-Target Attractor Control, that corresponds with the problem of identifying cellular reprogramming strategies was formulated. The pbn-STAC framework was shown to be able to find optimal and sub-optimal control strategies between source and target attractors for BN and PBN models of sizes up to 33 nodes.
@@
In this study, we build on the initial idea of [7] to devise a general and scalable approach capable of handling very large BN models of biological networks. To this end, we introduce the Graph-based Attractor-Target Control Algorithm (GATTACA) computational framework to solve a novel, general target-control problem for BN models of biological networks under the asynchronous update mode. Our new control problem is tailored for cellular reprogramming in the sense that control actions are allowed to take place only in attractor states as they correspond to observable phenotypic cellular states understood as discrete, mutually exclusive, stable, and observable cell phenotypes or cell fates that can undergo transitions from one to another under the influence of regulatory inputs. This approach addresses the constraints of wet-lab biology, where it is very difficult to experimentally determine a transient state of a cell, without destroying it, for the application of a control intervention in a real-world scenario. In contrast, stable cell phenotypes, which correspond to the attractor states of a dynamical system [8], are expected to be more easily, optically observable realisations of complex interactions between the genome, epigenome, and local environment. To reflect these constraints, we restrict control actions to be applied only in observable phenotypic cellular states, which are represented by attractors of a BN model.
@@
The core of our GATTACA framework is a Graph Neural Network (GNN)-based Q-learning agent which learns minimal control strategies that drive the system from a given source attractor state to one aligned with a given target configuration, i.e. an attractor with user-specified values for a subset of nodes. The structure of the biological network is encoded directly into the GNN, allowing the model to exploit topological relationships between the network components (nodes) during training. Furthermore, to facilitate effective identification of attractors in large BN models of complex networks, we consider the notion of pseudo-attractor states introduced in [7] and we enhance the pseudo-attractor states identification procedure proposed therein.
@@ II Related works
The application of reinforcement learning to control GRNs was pioneered in [9]. The study focused on avoiding undesirable states in terms of steady-state probabilities of PBNs. The authors treated time-series gene expression samples as a sequence of experience tuples and used a batch version of Q-Learning to approximate the optimal policy. Later, to relax the temporal condition, the BOAFQI-Sarsa method was devised in [10]. A further development was mSFQI method proposed in [11] for control based on the probabilities of gene activity profiles.
@@
Rule-based reinforcement learning in the form of an eXtended Classifier System (XCS) was devised in [12] for the control of random BNs. Subsequently, it was applied in [13] to a 11-node yeast cell cycle network.
@@
Scalability is the main concern when devising new methods for control of BNs. This issue can be addressed by devising new methods based on deep reinforcement learning (DRL) techniques, which have proven highly successful in decision problems characterized by large state-action spaces, see, e.g. [6]. To this end, the idea of combining Q-learning with deep learning was conceived. The first to apply this idea in the context of BN control were Papagiannis et al. in [14], where a control method for BNs based on deep Q-Learning with experience reply, namely Double Deep Q-Network (DDQN) with Proritised Experience Replay (PER), was proposed. It was applied in [15] to control a 20-node synthetic PBN and a 7-node real-world PBN model of melanoma. The same approach was used by Acernese et al. in [16] for the control of synchronous PBNs with the goal of driving the networks from a particular state towards a more desirable one. Next, in [17], the authors proposed a Q-learning based method to address the feedback stabilisation problem of probabilistic Boolean control networks (PBCNs) with the aim of stabilising the system at a given equilibrium point and applied it to a small 9-node model of apoptosis.
@@
To address the control problem on large networks, Moschoyiannis et al. in [18] applied DDQN with PER to control a large synchronous melanoma PBN model of 200 nodes. They limited the control actions to flipping a single specific gene, namely pirin, which had already been identified as relevant for the control of melanoma PBN models in previous studies. Furthermore, the target set of states consisted of exactly half of the entire state space, that is, the objective was to reach a subset of network states where the WNT5A gene is OFF.
@@
Finally, Mizera et al. in [7] leveraged DRL to develop pbn-STAC, a computational framework that solves the source-target attractor control problem of driving the network dynamics from a given source to a specified target attractor of a BN or PBN model under the asynchronous update mode. The control problem is tailored for cellular reprogramming in the sense that control actions are allowed to take place only in intermediate attractor states to reflect the constraints of wet-lab biology. The approach was successful in controlling BN and PBN models of up to 33-nodes.
@@ III-A Boolean networks
We formally define a Boolean network $B$ as a discrete dynamical system represented by a pair $B=(V,\mathcal{F})$, where $V=\{v_{1},v_{2},\ldots,v_{n}\}$ is a set of binary-valued nodes (a.k.a genes) and =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{n}\}$ is a set of predictor functions. Each predictor function $f_{i}\in\mathcal{F}$ is a Boolean function that encodes the state transition rule of node $v_{i}$. It is defined with respect to a subset of $V$ denoted $\textrm{Pa}(v_{i})\subseteq V$ and referred to as the parent nodes (regulators) of $v_{i}$ in $B$. The predictor function $f_{i}$ determines the state of $v_{i}$ at a discrete time point $t+1$ based on the states of the parent nodes at time point $t$.
@@
structure graph of a BN is a directed graph $G=(V,E)$, where $V$ is the set of nodes of the BN representing components of the system and $E$ is a set of edges defined as $E=\{(v,u)\in V\times V\mid v\in\textrm{Pa}(u)\}$. The structure graph represents the individual direct regulatory relationships between the components of the system encoded by the respective Boolean functions.
@@
The system dynamics evolves iteratively, with the states of the nodes being updated in accordance with their respective predictor functions. Two update modes are usually considered: synchronous, where all nodes are updated at once, and asynchronous, where one randomly chosen node is updated at a time. As in [7], also in this study we focus on the asynchronous mode, which is considered more appropriate for the modelling of biological networks [19, 20], yet more challenging from a computational point of view.
state $x$ of a Boolean network $B=(V,\mathcal{F})$ is a binary vector $x=(x_{1},x_{2},\ldots,x_{n})$, where $n=|V|$ is the number of nodes in the network and each $x_{i}\in\{0,1\}$ represents the state of node $v_{i}$.
@@
An associated state transition graph (STG) of $B$ under the asynchronous update mode is a directed graph that represents the possible state transitions of $B$. Nodes of the STG correspond to states of $B$, and there exists a directed edge from state $s$ to state $s^{\prime}$ in the STG if and only if there exists a predictor function in $\mathcal{F}$ such that applying it to state $s$ results in state $s^{\prime}$. Formally, $\textrm{STG}(B)=(S,\rightarrow)$, where $S=\{0,1\}^{|V|}$ is the set of all possible states of $B$, and $\rightarrow$ is the set of directed edges representing state transitions, i.e. =$\rightarrow=\{(s,s^{\prime})\in S\times S\mid\exists f_{i}\in\mathcal{F}\ f_{i}(s)=s^{\prime}\}$.
@@
An attractor of a BN is a bottom strongly connected component (BSCC) in the STG of the network. A fixed-point attractor and a multi-state attractor are BSCCs consisting of a single state and more than one state, respectively.
@@ III-B Pseudo-attractors
It was shown that obtaining the family of all attractors of a BN is NP-hard by itself [21]. This is especially prohibitive in the case of cellular reprogramming considerations, where ultimately large models need to be analysed. However, from a practical point of view in the context of cellular reprogramming, only the frequently revisited states of an attractor are the relevant ones since they correspond to phenotypical cellular states that are observable in the lab. This makes these states ‘recognisable’ for the application of reprogramming interventions in practice in accordance with the control strategy obtained with our framework. With this in mind, the notions of a pseudo-attractor and a pseudo-attractor state were introduced in [7]. Since in this study our focus is solely on BN models under the asynchronous update mode, herein we provide a variant of the definition explicitly referring to asynchronous BNs.
@@
In [7] it was proved that for any BN attractor there exists a non-empty pseudo-attractor as stated by the following two theorems.

### v3

@@ I Introduction
Complex diseases remain a major challenge for biomedical research due to the intricate interactions between genes forming gene regulatory networks (GRNs). These networks exhibit complex dynamics emerging from the structure of the network: the state of the genome evolves in time, finally reaching some stable gene expression profiles characterising cell types or, more generally, stable cellular functional states [HEBI05].
@@
In some cases, when a GRN is perturbed, it can reach an “unhealthy” state, that is otherwise inaccessible [Barabasi2011NetMed]. An artificial change of the cell fate, referred to as cellular reprogramming, aims to solve this issue by steering the GRN back to a “healthy” state. However, identifying effective reprogramming targets and intervention strategies through wet-lab experiments alone is prohibitively expensive and time-consuming. This motivates the development of in-silico methods.
@@
Capturing emergent properties of GRN dynamics requires the consideration of the system as a whole. To this end, GRNs can be holistically modelled with the formalism of Boolean networks (BNs) originally introduced by Kauffman in [KS69]. Its simplicity enables tractable simulations and analysis of non-linear biological dynamics while retaining essential, emergent complex system behaviour. Cellular reprogramming can be formulated as a control problem within the framework of BNs.
@@
To support cellular reprogramming, control strategies must be computed for large-scale BNs, taking into account both the structure and dynamics of the network [GR16]. Scalability is the main concern when devising new computational methods for control of asynchronous BNs. Existing state-of-the-art methods and tools, such as CABEAN [SP21], provide exact algorithms for attractor-based control. Nevertheless, they are limited in scalability because of the infamous “curse of dimensionality” problem, i.e. the size of the state space grows exponentially with the number of genes considered, which is handled by additional assumptions about network decomposability.
@@
Deep reinforcement learning (DRL) has proven highly successful in decision problems characterised by very large state-action spaces [MuZero20]. Therefore, it seems a promising alternative to traditional methods for controlling GRNs. In [ours_tcs], pbn-STAC – a DRL-based tool for control of asynchronous BNs and probabilistic Boolean networks (PBNs) was introduced, which was a result of exploration into the potential of DRL for developing scalable control methods for both BN and PBN models of GRNs in the context of cellular reprogramming. Therein, notions of a pseudo-attractor and a pseudo-attractor state were introduced, a pseudo-attractor states identification procedure (PASIP) was devised, and a novel control problem, i.e. the Source-Target Attractor Control, that corresponds with the problem of identifying cellular reprogramming strategies was formulated. The pbn-STAC framework was shown to be able to find optimal and sub-optimal control strategies between source and target attractors for BN and PBN models of sizes up to 33 nodes.
@@
In this study, we build on the initial idea of [ours_tcs] to devise a general and scalable approach capable of handling very large BN models of biological networks. To this end, we introduce the Graph-based Attractor-Target Control Algorithm (GATTACA) computational framework to solve a novel, general target-control problem for BN models of biological networks under the asynchronous update mode. Our new control problem is tailored for cellular reprogramming in the sense that control actions are allowed to take place only in attractor states as they correspond to observable phenotypic cellular states understood as discrete, mutually exclusive, stable, and observable cell phenotypes or cell fates that can undergo transitions from one to another under the influence of regulatory inputs. This approach addresses the constraints of wet-lab biology, where it is very difficult to experimentally determine a transient state of a cell, without destroying it, for the application of a control intervention in a real-world scenario. In contrast, stable cell phenotypes, which correspond to the attractor states of a dynamical system [Grunt25], are expected to be more easily, optically observable realisations of complex interactions between the genome, epigenome, and local environment. To reflect these constraints, we restrict control actions to be applied only in observable phenotypic cellular states, which are represented by attractors of a BN model.
@@
The core of our GATTACA framework is a Graph Neural Network (GNN)-based Q-learning agent which learns minimal control strategies that drive the system from a given source attractor state to one aligned with a given target configuration, i.e. an attractor with user-specified values for a subset of nodes. The structure of the biological network is encoded directly into the GNN, allowing the model to exploit topological relationships between the network components (nodes) during training. Furthermore, to facilitate effective identification of attractors in large BN models of complex networks, we consider the notion of pseudo-attractor states introduced in [ours_tcs] and we enhance the pseudo-attractor states identification procedure proposed therein.
@@ II Related works
The application of reinforcement learning to control GRNs was pioneered in [SPA13]. The study focused on avoiding undesirable states in terms of steady-state probabilities of PBNs. The authors treated time-series gene expression samples as a sequence of experience tuples and used a batch version of Q-Learning to approximate the optimal policy. Later, to relax the temporal condition, the BOAFQI-Sarsa method was devised in [NCB18]. A further development was mSFQI method proposed in [NBC20] for control based on the probabilities of gene activity profiles.
@@
Rule-based reinforcement learning in the form of an eXtended Classifier System (XCS) was devised in [Karlsen2018] for the control of random BNs. Subsequently, it was applied in [Karlsen2019] to a 11-node yeast cell cycle network.
@@
Scalability is the main concern when devising new methods for control of BNs. This issue can be addressed by devising new methods based on deep reinforcement learning (DRL) techniques, which have proven highly successful in decision problems characterized by large state-action spaces, see, e.g. [MuZero20]. To this end, the idea of combining Q-learning with deep learning was conceived. The first to apply this idea in the context of BN control were Papagiannis et al. in [Papagiannis2019], where a control method for BNs based on deep Q-Learning with experience reply, namely Double Deep Q-Network (DDQN) with Proritised Experience Replay (PER), was proposed. It was applied in [Papagiannis2021] to control a 20-node synthetic PBN and a 7-node real-world PBN model of melanoma. The same approach was used by Acernese et al. in [Acernese2020] for the control of synchronous PBNs with the goal of driving the networks from a particular state towards a more desirable one. Next, in [Acernese2021], the authors proposed a Q-learning based method to address the feedback stabilisation problem of probabilistic Boolean control networks (PBCNs) with the aim of stabilising the system at a given equilibrium point and applied it to a small 9-node model of apoptosis.
@@
To address the control problem on large networks, Moschoyiannis et al. in [Moschoyiannis2023] applied DDQN with PER to control a large synchronous melanoma PBN model of 200 nodes. They limited the control actions to flipping a single specific gene, namely pirin, which had already been identified as relevant for the control of melanoma PBN models in previous studies. Furthermore, the target set of states consisted of exactly half of the entire state space, that is, the objective was to reach a subset of network states where the WNT5A gene is OFF.
@@
Finally, Mizera et al. in [ours_tcs] leveraged DRL to develop pbn-STAC, a computational framework that solves the source-target attractor control problem of driving the network dynamics from a given source to a specified target attractor of a BN or PBN model under the asynchronous update mode. The control problem is tailored for cellular reprogramming in the sense that control actions are allowed to take place only in intermediate attractor states to reflect the constraints of wet-lab biology. The approach was successful in controlling BN and PBN models of up to 33-nodes.
@@ III-A Boolean networks
Boolean function $f$ is a function of $n\in\mathbb{N}$ Boolean variables, $f:\{0,1\}^{k}\to\{0,1\}$, whose arguments and output take values in the binary set $\{0,1\}$. A variable $x_{l}$ is an essential variable of a Boolean function $f$ if and only if $f(x_{1},\ldots,x_{l-1},0,x_{l+1},\ldots,x_{n})\neq f(x_{1},\ldots,x_{l-1},1,x_{l+1},\ldots,x_{n})$ for some fixed values of $x_{1},\ldots,x_{l-1},x_{l+1},\ldots,x_{n}$. A Boolean network $B$ is a discrete dynamical system defined as a tuple $B=(\boldsymbol{x},\boldsymbol{f}),$ where $\boldsymbol{x}=(x_{1},\allowbreak x_{2},\ldots,x_{n})$ is a vector of binary-valued variables (also referred to as nodes or genes) and $\boldsymbol{f}=(f_{1},\allowbreak f_{2},\ldots,f_{n})$ is a vector of Boolean functions. Each $x_{i}of\textbf{x}$ is associated with a predictor function (or simply a predictor) $f_{i}(x_{i_{1}},x_{i_{2}},\ldots,x_{i_{l(i)}}),$ which is the $i$-th component of $\boldsymbol{f}$ and is defined with respect to $1\leq l(i)\leq n$ distinct variables from $\boldsymbol{x}$. The set of variables $\{x_{i_{1}},x_{i_{2}},\ldots,x_{i_{l(i)}}\}$ is referred to as the set of parent nodes of $x_{i}$, denoted $\textrm{Pa}(x_{i})$. We will assume that parent nodes $\textrm{Pa}(x_{i})$ correspond to the set of essential variables of $f_{i}$. The predictor function $f_{i}$ defines the evolution of variable $x_{i}$ in discrete time steps.
@@
The structure graph of $B=(\textbf{x},\textbf{f})$ is the directed graph $SG_{\textrm{B}}=(V,E)$, where $V$ is the set of nodes consisting of the variables of x and $E=\{(x_{j},x_{i})\in V\times V\mid x_{j}\in\textrm{Pa}(x_{i})\}$. The structure graph represents the individual direct regulatory relationships between the components of the system encoded by the respective Boolean functions.
@@
The system dynamics evolves iteratively, with the states of the nodes being updated in accordance with their respective predictor functions. Two update modes are usually considered: synchronous, where all nodes are updated at once, and asynchronous, where one randomly chosen node is updated at a time. As in [ours_tcs], also in this study we focus on the asynchronous mode, which is considered more appropriate for the modelling of biological networks [CFM20, ZYLW+13], yet more challenging from a computational point of view.
@@
The state space of a BN with $n$ nodes is $S=\{0,1\}^{n}$. At each time point $t$, the state of the network is defined by the vector $\boldsymbol{x}(t)=(x_{1}(t),x_{2}(t),\ldots,x_{n}(t))$, where $x_{i}(t)$ is the value of variable $x_{i}$ at time $t$, i.e. $x_{i}(t)\in\{0,1\}$. Under the asynchronous update mode, a single randomly selected node is updated at each time step according to its corresponding predictor function. The associated asynchronous state transition graph (STG) of $B$, denoted $\mathrm{STG}(B)$, is a directed graph that represents the possible state transitions of $B$ under the asynchronous update mode. Formally, $\mathrm{STG}(B)=(S,\rightarrow)$, where $S=\{0,1\}^{n}$ is the set of all possible states and $\rightarrow$ is the set of directed edges such that a directed edge from $\mathbf{s}=(s_{1},s_{2},\ldots,s_{n})\in S$ to $\mathbf{s^{\prime}}=(s_{1}^{\prime},s_{2}^{\prime},\ldots,s_{n}^{\prime})\in S$, denoted $\mathbf{s}\rightarrow\mathbf{s^{\prime}}$, is in $\rightarrow$ if and only if $\mathbf{s^{\prime}}$ can be obtained from $\mathbf{s}$ by a single asynchronous update, i.e. there exists $i=1,2,\ldots,n$, such that
$\begin{array}[]{lcl}\left(s_{i}^{\prime}=f_{i}(s_{i_{1}},s_{i_{2}},\cdots,s_{i_{l(i)}})\right)\wedge\left(\bigwedge_{j\neq i}\ (s_{j}^{\prime}=s_{j})\right).\end{array}$
For a BN $B$ under the asynchronous update mode, a path from a state $\mathbf{s}$ to a state $\mathbf{s^{\prime}}$ is a (possibly empty) sequence of transitions from $\mathbf{s}$ to $\mathbf{s^{\prime}}$ in $\mathrm{STG}(B)$. A path from a state $\mathbf{s}$ to a subset $S^{\prime}$ of $S$ is a path from $\mathbf{s}$ to any state $\mathbf{s^{\prime}}\in S^{\prime}$. For any state $\mathbf{s}\in S$, let $\mathsf{pre}(\mathbf{s})=\{\mathbf{s^{\prime}}\in S\mid\mathbf{s^{\prime}}\rightarrow\mathbf{s}\}$ and let $\mathsf{post}(\mathbf{s})=\{\mathbf{s^{\prime}}\in S\mid\mathbf{s}\rightarrow\mathbf{s^{\prime}}\}$. $\mathsf{pre}(\mathbf{s})$ contains all the states that can reach $\mathbf{s}$ by performing a single transition in $\mathrm{STG}(B)$ and $\mathsf{post}(\mathbf{s})$ contains all the states that can be reached from $\mathbf{s}$ by a single transition in $\mathrm{STG}(B)$. The definition of $\mathsf{post}$ can be lifted to a subset $S^{\prime}\subseteq S$ as:
$\mathsf{post}(S^{\prime})=\bigcup_{\mathbf{s\in S^{\prime}}}(\mathsf{post}(\mathbf{s})).$
For a state $\mathbf{s}\in S$, $\mathsf{reach}(\mathbf{s})$ denotes the set of states $\mathbf{s^{\prime}}$ such that there is a path from $\mathbf{s}$ to $\mathbf{s^{\prime}}$ in $\mathrm{STG}(B)$ and can be defined as the transitive closure of the $\mathsf{post}$ operation, i.e. we define:
$\mathsf{post}^{0}(\mathbf{s})$ $=\{\mathbf{s}\},$
$\mathsf{post}^{i+1}(\mathbf{s})$ $=\mathsf{post}(\mathsf{post}^{i}(\mathbf{s})),$
and
$\mathsf{reach}(\mathbf{s})=\bigcup_{i=0}^{\infty}(\mathsf{post}^{i}(\mathbf{s})).$
Thus, $\mathsf{reach}(\mathbf{s})$ is the smallest subset of states in $S$ such that $\mathbf{s}\in\mathsf{reach}(\mathbf{s})$ and $\mathsf{post}(\mathsf{reach}(\mathbf{s}))\subseteq\mathsf{reach}(\mathbf{s})$.
@@
An attractor $A$ of $\mathrm{STG}(B)$ is a subset of states of $S$ such that for every $\mathbf{s}\in A$, $\mathsf{reach}(\mathbf{s})=A$. It is a bottom strongly connected component (BSCC) in the STG of the network. A fixed-point attractor and a multi-state attractor are BSCCs consisting of a single state and more than one state, respectively.
@@ III-B Pseudo-attractors
It was shown that obtaining the family of all attractors of a BN is NP-hard by itself [Akutsu2003]. This is especially prohibitive in the case of cellular reprogramming considerations, where ultimately large models need to be analysed. However, from a practical point of view in the context of cellular reprogramming, only the frequently revisited states of an attractor are the relevant ones since they correspond to phenotypical cellular states that are observable in the lab. This makes these states ‘recognisable’ for the application of reprogramming interventions in practice in accordance with the control strategy obtained with our framework. With this in mind, the notions of a pseudo-attractor and a pseudo-attractor state were introduced in [ours_tcs]. Since in this study our focus is solely on BN models under the asynchronous update mode, herein we provide a variant of the definition explicitly referring to asynchronous BNs.
@@
In [ours_tcs] it was proved that for any BN attractor there exists a non-empty pseudo-attractor as stated by the following two theorems.